OpenStudy (anonymous):

Can someone help me with some Algebra 2 questions?

OpenStudy (anonymous):

\[\sqrt{2x-3}=5\]

OpenStudy (anonymous):

Solve for x.

OpenStudy (amistre64):

the square root of what number is 5?

OpenStudy (anonymous):

2?

OpenStudy (amistre64):

sqrt(2) = 5 ?? no

OpenStudy (anonymous):

Well, I don't know. I'm stupid..seriously when it comes to algebra

OpenStudy (amistre64):

this isnt algebra .... there is an algebraic want to approach it, but it would also be useful to know what a square and square root tell us

OpenStudy (amistre64):

think about what the question is asking for: the square root of some number is 5, therefore some number is equal to 5^2

OpenStudy (amistre64):

if we square both sides, we can unwrap the square root \[(\sqrt{...})=k\] \[(\sqrt{...})^2=k^2\] \[(...)=k^2\]

OpenStudy (anonymous):

i don't know.............

OpenStudy (anonymous):

i'm so frustrated right now it's not even funny

OpenStudy (amistre64):

5^2 is notation that tells us how many times to multiply a number with itself .... 5^2 means; 5*5

OpenStudy (amistre64):

if \(5^2 = n\), then \(\sqrt n = 5\)

OpenStudy (anonymous):

this is all foreign language..... you don't have to help me if you don't want to i'm obviously an idiot.

OpenStudy (amistre64):

lets assume for the moment that 5^2 = 25

OpenStudy (amistre64):

then the sqrt(25) = 5 they give us; sqrt(2x-3) = 5 this tells us that 2x-3 has to equal 25 2x-3 = 25, should be readily solvable for x

OpenStudy (amistre64):

\[\sqrt{2x-3}=5~:~square ~both~sides\] \[(\sqrt{2x-3})^2=5^2\] \[2x-3=25\] that should be the hardest part about this ... i trust you can work the rest of it out

OpenStudy (anonymous):

No i can't like i said i'm terrible at this and i'm probably going to fail, but thanks anyway

OpenStudy (anonymous):

@KelseyKaup123 , amistre did all the hard work, not it is just simple Algebra or Pre-Algebra even! You can do it!! you want to get "x" all by itself ... so you need to shift things around to do that first take that "-3" and switch it to the other by add it +3 to both sides, which will then give you 2x = 25+3 Can you solve for that?